# A Midsummer Meme’s Dream: Investigating Market Manipulations in the Meme Coin Ecosystem

### v1

@@ 5.3.1. LPI Detection
The second phase implements a more stringent verification step to minimize false positives. Using the Dune Analytics API, we retrieve granular transaction-level data for each flagged event and introduce two additional constraints. First, buy volume must account for at least 90% of the total daily trading volume, indicating sustained buying pressure rather than balanced market activity. Second, the number of unique trading entities must not exceed 10, implying coordinated or strategic behavior rather than organic participation by a broad user base. This methodology detects 72 price manipulation events affecting 40 distinct meme coins, 12 on Solana, 11 on BSC, 10 on Ethereum, and 7 on Base.
5.3.2. LPI Analysis
@@
Figure 5. Relationship between price increase percentage and buy volume for LPI activities across different blockchains.
[…]
The figure shows a scatter plot showing the relationship between price increase percentage (x-axis) and buy volume (y-axis) for LPI activities across different blockchains.
High Price Impact with Minimal Investment. Fig. 5 shows the relationship between buy volume and price impact across the detected events revealing that manipulators achieved substantial price increases with minimal investment. Indeed, the median buy volume is merely \$$\$54$, with only $43.06\%$ (31) of events requiring more than \$$\$100$ in capital. Ethereum stands apart from other blockchains with its \$560 median, while others show median values below \$5. Despite these modest volumes, the effects on the prices are dramatic: $50\%$ of manipulations generate price increases exceeding $500\%$, while $36.11\%$ produce gains surpassing $1,000\%$. In particular, the most extreme price distortions (exceeding $10,000\%$) occur with the smallest capital inputs, highlighting the vulnerability of low liquidity pools. The manipulation effectiveness varied significantly across blockchains: Base demonstrated the highest median price increase at $4\times 10^{8}\%$, followed by Solana ($661\%$), BSC ($504\%$), and Ethereum ($214\%$).
@@
Targets Are Old Tokens. Unlike other manipulation techniques that frequently target newly launched tokens (Cernera et al., 2023a, b), LPI occurs predominantly at the more mature stages of the lifecycle of a token. Only 2 meme coins experience manipulation within their first month after launch, and only 9 tokens are manipulated within their first 90 days. The median time to manipulation is 209 days after token creation, suggesting that manipulators specifically seek out older tokens that never gained significant traction. These tokens present ideal targets as they appear more legitimate due to their longer existence while still maintaining the low liquidity pools necessary for effective price manipulation. This temporal pattern is particularly different on Solana, where tokens are targeted much earlier, with a median of just 115 days, compared to tokens on other blockchains, which have a median of over 163 days.
@@
5.4. Legitimate vs. Manipulated Growth
@@
The analysis of high-performing tokens shows that the vast majority display signs of market manipulation or conditions that facilitate it. By combining results from wash trading (282 tokens), LPI strategies (40), and strong anomaly indicators (412), we identify 584 tokens (82.6% of all high-performing tokens) that exhibit at least one form of suspicious activity. In particular, 117 tokens (20.03%) show wash trading and strong anomaly indicators, with 114 tokens exhibiting extreme ownership concentration due to the top 10 accounts holding an average of 73.09% of the supply. Ownership concentration is even more pronounced in cases related to LPI, where of the 40 involved tokens, 32 show clear signs of concentrated ownership. This includes 27 tokens where the top 10 holders control an average of 80.62% of the supply, and two cases where all the supply is acquired through bundle buy transactions, likely to obscure actual ownership. Furthermore, we observe 11 tokens involved in both wash trading and LPI activities. This combination of tactics suggests coordinated efforts to inflate trading volume and price, creating a convincing illusion of strong market demand and growth.
@@
This analysis leaves only 123 tokens that show no signs of market manipulation or suspicious ownership. In terms of blockchain distribution, these successful tokens are primarily on Solana (82 tokens), followed by BSC (22), Ethereum (12), and Base (7). When normalized by the total number of tokens on each platform for which we have price data (see Section 4.2.1), Base has the highest success rate (with 1.14%), followed by Solana (0.79%), Ethereum (0.62%), and BSC with the lowest success rate (0.30%).
6. Profit Extraction Manipulations
@@
Following our discovery that 82.6% of high-performing meme coins exhibited manipulative practices (Sec. 5.4), this section explores whether such artificial growth is part of a calculated strategy to extract profits at the expense of investors. To track this progression from manipulation to extraction, we extended our analysis with OHLCV data from mid-January through mid-March 2025. This additional window allows us to observe whether tokens that experienced price increases are later involved in exit strategies.
6.1. Strategic Exit via Rug Pulls
@@
Rug pulls represent a severe form of exit scam where developers abandon their project after generating significant initial investment, typically by removing all liquidity from trading pools (Zhou et al., 2024). To detect rug pulls among high-performing tokens, we used two strict conservative criteria: (1) a single-day price drop of over 99%, signaling that the token has effectively lost all value, a decline too extreme to be explained by normal market fluctuations, and (2) a sustained drop in trading volume of more than 99% in the following days, indicating that liquidity was removed and investors could no longer trade the token. These conditions confirm that the token has been abandoned and has not experienced a temporary market dip.
Applying this approach to our updated OHLCV data, we detect two tokens subjected to apparent rug pulls: one on Solana and the other on Ethereum. The timing of these events reveals calculated planning, with developers executing rug pulls 248 and 451 days after token creation, respectively. This suggests a strategy of establishing credibility over an extended period before executing the exit. Moreover, both tokens have previously undergone other market manipulations to generate artificial interest, one involved in wash trading and the other in price pumping, indicating a progressive manipulation strategy culminating in the final exit scam.
6.2. Orchestrated Pump and Dump Operations
@@
Pump and dumps are another common manipulation in which the price of a token is artificially inflated through coordinated buying, followed by rapid selling that crashes the price and leaves late investors with substantial losses (Xu and Livshits, 2019). Our methodology for detecting pump and dump operations is the following:
@@
1. Price peaks detection: We identify price peaks demonstrating overbought conditions with a Relative Strength Index exceeding 80, indicating excessive buying momentum (Ţăran-Moroşan, 2011). For each detected peak, we examine the preceding 24-hour window to locate the local minimum price point as the potential manipulation starting point.
@@
To minimize false positives, we implemented more conservative thresholds than previous research by Kamps et al. (Kamps and Kleinberg, 2018), who employed a 10% price increase threshold with a 400% volume surge requirement. This approach detects 91 pump and dump operations affecting 60 different tokens, 54 on Solana, 3 on BSC, 2 on Base, and 1 on Ethereum.
@@
Figure 6. Relationship between price increase during pump and decrease during dump across different blockchains.
[…]
Scatterplot showing the magnitude of price increases during pump phases (x-axis, logarithmic scale) and subsequent price decreases during dump phases (y-axis) for each pump and dump operation, color-coded by blockchain.
@@
As shown in Fig. 6, the pump phases exhibit significant variability in magnitude. The median price increase is 143.95%, with 63.74% of operations (58) featuring increases of at least 100% and 7 operations (7.69%) with increases exceeding 1,000%. The subsequent dump phases show correspondingly severe price collapses, with a median decrease of 55.82%. In 59.34% of cases (54), tokens lost at least 50% of their value, while 17.58% (16) experienced severe dumps exceeding 80% price decline. Lastly, concerning the time elapsed between the creation date of meme coins and their first pump and dump, we have that, on average, passes 206 days, with a median of 146 days and a minimum of 49 days.
@@
The 60 affected tokens reveal a concerning pattern: 61.67% (37 tokens) were previously involved in artificial growth strategies, 34 in wash trading, and 3 LPI. If we also consider the rug pulled tokens, this ratio increases to 62.9%. This finding strongly suggests that manipulations examined in Sec. 5 often function as preliminary steps in a broader scheme designed to attract investors before profit extraction through pump and dumps or rug pulls.
6.3. Case Study: Delisted Tokens
@@
The removal (delisting) of tokens from major aggregators like CoinMarketCap and CoinGecko can signal either project abandonment or suspicious market behavior. To better understand manipulation tactics across this ecosystem lifecycle, we analyze tokens that were once listed, having met the platforms’ stringent listing criteria, but were later removed. These tokens represent a particularly interesting case study, as they gained the credibility and visibility associated with major platform listings before their removal. Our investigation focuses on two key questions: whether these tokens employed artificial growth techniques to satisfy initial listing criteria, and whether they subsequently exploited their enhanced visibility through market manipulations.
We examine 124 meme coins delisted from CoinMarketCap and CoinGecko between October 2024 and March 2025 and report the results in Tab. 5. Among these, 84 tokens (67.74%) showed signs of artificial growth manipulation, with 70 tokens exhibiting wash trading behavior, while 27 tokens employed LPI tactics. Additionally, 30 tokens (24.19%) show evidence of profit extraction manipulation after listing: 23 tokens present pump and dump patterns, while 7 are involved in rug pulls. Of these, 26 tokens (86.67%) had previously employed artificial growth tactics. Among the pump and dump tokens, 14 showed evidence of wash trading and 5 LPI. Likewise, all rug-pulled tokens showed signs of prior manipulation, with 5 involving wash trading and 6 LPI (4 of which exhibited both).
Table 5. Manipulation strategies in delisted meme coins
@@
7. Discussion and Ethical Considerations
Our study reveals that market manipulations are widespread among high-performing meme coins, posing significant risks to unsuspecting investors. DEX aggregators are failing to protect and inform users adequately. While services like DexScreener and Gecko Terminal offer security metrics, they place critical warnings at the bottom of pages, where they are easily missed by users making quick investment decisions. Moreover, technical metrics like token concentration percentages are difficult for average users to understand. Even more concerning is that these platforms do not provide alerts about specific manipulation tactics such as wash trading or LPI, which are often used to simulate fake interest and inflate token values. To improve user protection, these platforms should redesign their interfaces to highlight risk indicators prominently at the top of token pages. These indicators should be clear, simple, and designed to help everyday users understand when a token displays signs of manipulation. This issue becomes even more pressing when considering that we observe recurring manipulation strategies on tokens previously listed on established aggregators such as CoinMarketCap and CoinGecko. These platforms, viewed as reliable due to their stricter listing requirements, inadvertently lend credibility to manipulated tokens. To better safeguard users, these aggregators must enhance their verification processes by introducing checks that detect artificial growth strategies, tactics we have identified as early steps in broader profit extraction schemes.
Concerning ethical considerations, all data examined in this work is publicly available on blockchains, and the addresses are pseudo-anonymous. Throughout our investigation, we made no attempts to correlate addresses with external information that could compromise user privacy. Consequently, according to our IRB’s policy, we did not need explicit authorization to perform our study.
8. Limitations
The nature of our cross-chain analysis, encompassing four different blockchains, made direct on-chain data extraction computationally prohibitive. This limitation is particularly pronounced for high-throughput blockchains like Solana, where complete transaction history analysis would require substantial infrastructure investment. To overcome these constraints while maintaining analytical rigor, we leveraged third-party APIs: Gecko Terminal for ecosystem-wide OHLCV data, and public nodes and Dune Analytics for transaction-level examination of high-performing tokens.
To prioritize precision over recall, our market manipulation detections employ deliberately conservative thresholds selected to identify clear cases where trading patterns significantly deviate from natural market behavior. For instance, a 500% volume surge with minimal price movement represents an anomaly difficult to attribute to legitimate trading. This conservative approach introduces a detection bias toward more extreme manipulation cases, potentially missing subtler distortions. However, this methodological choice significantly increases our confidence in the validity of the detected manipulations and ensures that our manipulation rates represent a reliable lower bound rather than speculative estimates.
9. Conclusion and Future Work

### v2

@@ 5.3.1 LPI Detection
The second phase implements a more stringent verification step to minimize false positives. Using the Dune Analytics API, we retrieve granular transaction-level data for each flagged event and introduce two additional constraints. First, buy volume must account for at least 90% of the total daily trading volume, indicating sustained buying pressure rather than balanced market activity. Second, the number of unique trading entities must not exceed 10, implying coordinated or strategic behavior rather than organic participation by a broad user base. This methodology detects 72 price manipulations affecting 40 distinct meme coins, 12 on Solana, 11 on BSC, 10 on Ethereum, and 7 on Base.
5.3.2 LPI Analysis
@@
Figure 5: Relationship between price increase percentage and buy volume for LPI activities across different blockchains.
High Price Impact with Minimal Investment. Fig. 5 shows the relationship between buy volume and price impact across the detected events revealing that manipulators achieved substantial price increases with minimal investment. Indeed, the median buy volume is merely $\mathdollar 54$, with only $43.06\%$ (31) of events requiring more than $\mathdollar 100$ in capital. Ethereum stands apart from other blockchains with its \$560 median, while others show median values below \$5. Despite these modest volumes, the effects on the prices are dramatic: $50\%$ of manipulations generate price increases exceeding $500\%$, while $36.11\%$ produce gains surpassing $1,000\%$. In particular, the most extreme price distortions (exceeding $10,000\%$) occur with the smallest capital inputs, highlighting the vulnerability of low liquidity pools. The manipulation effectiveness varied significantly across blockchains: Base demonstrated the highest median price increase at $4\times 10^{8}\%$, followed by Solana ($661\%$), BSC ($504\%$), and Ethereum ($214\%$).
@@
Targets Are Old Tokens. Unlike other manipulations that target newly launched tokens [cernera2023token, cernera2023ready], LPI occurs predominantly at the more mature stages of the lifecycle of a token. Only 2 meme coins experienced manipulation within their first month, and only 9 within their first 90 days. The median time to manipulation is 209 days after token creation, suggesting that manipulators specifically seek out older tokens that never gained significant traction. These tokens present ideal targets as they appear more legitimate due to their longer existence while still maintaining the low liquidity pools necessary for effective price manipulation. This temporal pattern is particularly different on Solana, where tokens are targeted much earlier, with a median of just 115 days, compared to tokens on other blockchains, which have a median of over 163 days.
@@
5.4 Legitimate vs. Manipulated Growth
@@
The analysis of high-performing tokens shows that the vast majority display signs of market manipulation or conditions that facilitate it. By combining results from wash trading (287 tokens), LPI strategies (40), and strong anomaly indicators (412), we identify 586 tokens (82.89% of all high-performing tokens) that exhibit at least one form of suspicious activity. In particular, 120 tokens (20.48%) show wash trading and strong anomaly indicators, with 95 tokens exhibiting extreme ownership concentration due to the top 10 accounts holding an average of 72.94% of the supply. Ownership concentration is even more pronounced in cases related to LPI, where of the 40 involved tokens, 32 show clear signs of concentrated ownership. This includes 27 tokens where the top 10 holders control an average of 80.62% of the supply, and two cases where all the supply is acquired through bundle buy transactions, likely to obscure actual ownership. Furthermore, we observe 12 tokens involved in both wash trading and LPI activities. This combination of tactics suggests coordinated efforts to inflate trading volume and price, creating a convincing illusion of strong market demand and growth.
@@
This analysis leaves only 121 tokens that show no signs of market manipulation or suspicious ownership. In terms of blockchain distribution, these successful tokens are primarily on Solana (80 tokens), followed by BSC (22), Ethereum (12), and Base (7). When normalized by the total number of tokens on each platform for which we have price data (see Section 4.2.1), Base has the highest success rate (with 1.14%), followed by Solana (0.77%), Ethereum (0.62%), and BSC with the lowest success rate (0.30%).
6 Profit Extraction Manipulations
@@
With 82.8% of high-performing meme coins showing manipulative practices (Sec. 5.4), this section explores whether such artificial growth is part of a calculated strategy to extract profits at the expense of investors. To track this progression from manipulation to extraction, we extended our analysis using OHLCV data with hourly granularity from mid-January through mid-March 2025.
6.1 Strategic Exit via Rug Pulls
@@
Rug pulls represent a severe form of exit scam where developers abandon their project after generating significant initial investment, typically by removing all liquidity from trading pools [zhou2024stop]. To detect rug pulls among high-performing tokens, we applied the labeling criteria and thresholds used in [mazorra2022not]: (1) a single-day price drop of over 99%, signaling that the token has effectively lost all value, a decline too extreme to be explained by normal market fluctuations, and (2) a sustained drop in trading volume of more than 99% in the following days, indicating that liquidity was removed and investors could no longer trade the token. These conditions confirm that the token has been abandoned and has not experienced a temporary market dip. Applying this approach to our updated OHLCV data, we detect two tokens subjected to apparent rug pulls: one on Solana and the other on Ethereum. In particular, both tokens have previously undergone other market manipulations to generate artificial interest, one involved in wash trading and the other in LPI, indicating a progressive manipulation strategy culminating in the final exit scam.
6.2 Orchestrated Pump and Dump Operations
@@
Pump and dumps are another common manipulation in which the price of a token is inflated through coordinated buying, followed by rapid selling that crashes the price and leaves late investors with substantial losses [xu2019anatomy]. Our methodology for detecting pump and dump operations is the following:
@@
1. Price peaks detection: We identify price peaks demonstrating overbought conditions with a Relative Strength Index exceeding 80, indicating excessive buying momentum [ctuaran2011relative]. For each detected peak, we examine the preceding 24-hour window to locate the local minimum price point as the potential manipulation starting point.
@@
We initially adopted the detection thresholds from [kamps2018moon], who defined a pump as a 10% price increase with a 400% volume surge. However, as these parameters yielded excessive false positives, due to the volatility of meme coins, we implemented more conservative thresholds to prioritize precision over recall. This approach detects 91 pump and dump operations affecting 60 different tokens, 54 on Solana, 3 on BSC, 2 on Base, and 1 on Ethereum.
@@
Figure 6: Relationship between price increase during pump and decrease during dump across different blockchains.
@@
As shown in Fig. 6, the pump phases exhibit significant variability in magnitude. The median price increase is 143.95%, with 63.74% of operations (58) featuring increases of at least 100% and 7 operations (7.69%) with increases exceeding 1,000%. The subsequent dump phases show correspondingly severe price collapses, with a median decrease of 55.82%. In 59.34% of cases (54), tokens lost at least 50% of their value, while 17.58% (16) experienced severe dumps exceeding 80% price decline.
@@
The 60 affected tokens reveal a concerning pattern: 61.67% (37 tokens) were previously involved in artificial growth strategies, 35 in wash trading, and 3 LPI. If we also consider the rug pulled tokens, this ratio increases to 62.9%. This finding suggests that manipulations examined in Sec. 5 often function as preliminary steps in a broader scheme designed to attract investors before performing pump and dumps or rug pulls.
6.3 Case Study: Delisted Tokens
@@
The removal (delisting) of tokens from major aggregators like CoinMarketCap and CoinGecko can signal either project abandonment or suspicious market behavior. To examine manipulations across this lifecycle, we analyze tokens that were once listed, having met the platforms’ listing criteria, but were later removed. These tokens represent an interesting case study, as they gained the credibility and visibility associated with major platform listings before their removal. We investigate whether these tokens used artificial growth to meet listing criteria and then exploited their increased visibility through market manipulation. We examine 124 meme coins delisted from CoinMarketCap and CoinGecko between October 2024 and March 2025 and report the results in Tab. 5. Among these, 84 tokens (67.74%) showed signs of artificial growth, with 70 tokens exhibiting wash trading behavior, while 27 tokens employed LPI tactics. Additionally, 30 tokens (24.19%) show evidence of profit extraction after listing: 23 were involved in pump and dumps and 7 in rug pulls. Of these, 26 tokens (86.67%) had previously employed artificial growth tactics. Among the pump and dump tokens, 14 showed evidence of wash trading and 5 LPI. Likewise, all rug-pulled tokens showed signs of prior manipulation, with 5 involving wash trading and 6 LPI (4 of which exhibited both).
Table 5: Manipulation strategies in delisted meme coins
@@
6.4 Economic Impact
To quantify the financial harm posed by these manipulations, we analyzed the profit and loss flows for tokens involved in pump and dumps and rug pulls. Tab. 6 summarizes the profits and losses across these events.
Table 6: Profit and Loss (PnL) for pump and dump (P&P) participants and rug pull victims. Avg/Op is the mean PnL per operation (in \$), Avg/Addr is the mean PnL per wallet (in \$).
#Addr. Tot. PnL Avg/Op Avg/Addr
P&D
Creators 10,614 2.06M 26,034 779
Early Gainers 3,245 3.93M 51,652 579
Late Gainers 4,709 1.79M 22,436 365
Total Profit 18,568 7.78M 93,701 470
Early Losers 2,726 -1.92M -25,647 -724
Late Losers 3,057 -1.35M -17,529 -263
Total Loss 5,783 -3.27M -40,411 -617
Rug Pull
Victims 11,368 -6.04M -671,666 379
Pump and Dump Impact. We categorize pump and dump participants into three groups: creators, those who deployed the token or received it through airdrops, early investors, who purchased before the manipulation, and late investors, who bought during the pump phase. Creators and early/late gainers extracted a combined \$7.78 million in profits. Conversely, investors who sold at a loss suffered \$3.27 million in realized losses. This figure is a conservative lower bound: It only accounts for realized losses, not considering the unrealized losses of victims still holding the now-devalued tokens. The profit distribution reveals a clear exploitation hierarchy. Token creators extracted \$2.06 million across 10,614 addresses, averaging \$779 per wallet. This unusually high address count likely reflects the use of multiple wallets to obscure true ownership. Early investors who successfully exited accumulated \$3.93 million across 3,245 addresses (\$579 per wallet), while late investors achieved \$1.79 million across 4,709 addresses (\$365 per wallet). In total, 18,568 addresses realized gains averaging \$470 per wallet. However, these gains came at the expense of 5,783 addresses that suffered realized losses averaging \$617 per wallet. Early investors who sold at a loss experienced \$1.92 million in total losses (\$724 per wallet), while late investors suffered \$1.35 million (\$263 per wallet).
Rug Pulls Impact. Unlike pump and dumps where we track only realized losses, rug pulls effectively trap all remaining holders, as liquidity is removed from the market. Thus, we identify victims as addresses that purchased tokens before the rug pull and were unable to sell. The total losses from rug pulls reached \$6.04 million, with an average of \$671,666 per event, and with the most destructive operation causing losses exceeding \$2.89 million. In particular, rug pulls affected a total of 11,368 addresses, with an average of 1,263 addresses per event and one operation alone impacting 4,380 wallets.
Overall, we document a total financial loss exceeding \$9.3 million, impacting over 17,000 victim addresses. This highlights that the harm from these manipulations is widely distributed among retail investors, while the profits remain concentrated in the hands of creators and insiders, who likely employ multiple addresses to conceal their ownership concentration (as shown in Sec. 5.1.1).
7 Discussion
Implications for Ecosystem Stakeholders. Our findings reveal critical deficiencies in how cryptocurrency platforms present market data and the inadequacy of existing protection mechanisms. Security metrics, such as token concentration, require specialized knowledge to interpret and are typically buried at the bottom of token pages. Most critically, platforms fail to alert users to artificial growth strategies, relying instead on volume and liquidity metrics that our results demonstrate can be easily manipulated. To maintain their role as reliable information sources, these platforms must implement two key reforms. First, risk indicators should be displayed prominently using accessible language rather than technical metrics. Second, established aggregators must expand verification beyond basic technical checks to include continuous monitoring for artificial growth, which our analysis shows frequently precedes exploitation schemes. Moreover, this work supports regulators enforcing frameworks like MiCA [mica] by providing a comprehensive analysis of the manipulations, identifying detection patterns, quantifying the economic prevalence, and revealing novel strategies, such as LPI.
Countermeasures. Exit schemes such as rug pulls and pump-and-dumps execute rapidly, leaving minimal reaction time. Nonetheless, our work reveals that 62.9% of tokens subjected to profit extraction (86.67% among delisted tokens) had previously undergone artificial growth. This temporal pattern enables a viable defensive strategy: Detecting wash trading and LPI activities establishes an actionable warning window before profit extraction occurs. In this study, we employed conservative thresholds to minimize false positives, meaning our results likely underestimate total manipulation volume. Operational systems must carefully calibrate the precision-recall trade-off: Aggressive detection may identify more manipulations but risks flagging legitimate tokens, thereby undermining user trust, while overly conservative detection may fail to protect investors. Beyond technical detection, the most robust countermeasure remains investor education. Informed users who can recognize manipulation mechanics are unlikely to provide the exit liquidity that enables these schemes.
8 Conclusion and Future Work